Determine the speed of a motorboat in still water and the speed of the river current, if it takes 3 hours to travel a distance of 45 miles upstream and 2 hours to travel 50 miles downstream.
Answer by scott8148(6628)

b - c = 45/3 = 15

b + c = 50/2 = 25

adding equations (eliminating c) ___ 2b = 40